### v1

@@ VI-B Impact of Spatial Correlation
VI-C Impact of the Number of Users
@@
VI-D Impact of the Number of Antennas per User
@@
VI-E Impact of the Coherence Block Length
In Fig. 6, we look at the impact of the coherence block length on the SE with the ZF combiner. We vary the number of user antennas as in Fig. 5 and will revisit the issue mentioned in the former subsection. We note that the number of pilots increases linearly with the number of user antennas. Having more antennas should not degrade the SE, but it does when using ZF with $\tau_{\mathrm{c}}=200$. The reason is that the SE during data transmission increases little compared to the growing pilot overhead. However, if we increase $\tau_{\mathrm{c}}$, the extra pilot overhead is less impactful, relatively speaking. We can see that when $\tau_{\mathrm{c}}=1000$, the SE increases with $M$ for the considered range of antennas. Hence, the benefit of having more user antennas is particularly large in low-mobility cases. From Fig. 5 and Fig. 6 we see that the MMSE combiner is a better option than the ZF combiner. However, if the ZF combiner is to be preferred, higher $\tau_{\mathrm{c}}$ should be preferred.
Figure 6: Achievable per user SE with ZF combiner as a function of the number of user antennas.
VI-F Comparison of the Capacity Bounds
@@
Fig. 7 shows the average SE obtained with Method 1 for $M=1$ antenna per user. This is the most common setup in previous works, including the textbook [4]. There is little difference between the curves since channel hardening occurs in this case, so one can omit downlink pilots and use the hardening bound without much performance loss.
Fig. 8 considers the same setup but with $M=2$. As discussed earlier in the paper, we cannot rely on channel hardening because only the diagonal entries of the effective channel harden, while the inter-stream interference in the off-diagonal entries does not vanish. Consequently, there is a huge gap between the perfect CSI and no CSI cases for $M>1$. However, using estimates of the downlink effective channels, as proposed in Theorem 1 leads to vastly higher performance than in the case of no CSI. The difference between the perfect CSI and the proposed method is due to two primary factors. One is the ability of the combiner filter to suppress interference and mitigate estimation error, which is inherently dependent on the effective channel estimates. Secondly, it is affected by the increased channel estimation overhead.
In [4], it is briefly mentioned that multi-antenna users can be treated as multiple single-antenna users placed close to each other. Here, if we assume that each serving AP transmits a separate stream to each antenna of user $k$, our model aligns with that assumption. In Fig. 9, we compare the average SE achieved by this sub-optimal method with our proposed Method 3. The results are shown as a function of the number of antennas on each AP, and we consider both the proposed bound and the hardening bound. We observe that the hardening bound is close to the proposed bound when using the sub-optimal method. Therefore, the system can operate without downlink pilots with this transmission method but rely on channel hardening, leading to less channel estimation overhead. However, we see that the SEs achieved by this sub-optimal transmission method are significantly lower than those of Methods 1-3. This is due to the loss of beamforming gain and increased inter-stream interference. These results tell us that we should not treat a multi-antenna user as multiple single-antenna users, and we need to do pilot-based downlink channel estimation in order to have high SE with multiple antenna users. We also observe that, although the structure of the precoder is block diagonal for Method 3, the hardening bound isn’t tight. We omitted the SE results for Method 2 as these are similar to Method 3.
@@
[…]
@@
VII Conclusions
In this work, we present a novel analysis of cell-free massive MIMO systems with multi-antenna users, where each user can receive multiple spatial streams. We analyzed the transmission of these streams from either multiple APs, which requires data sharing, or from separate APs, and compared CSI-sharing and no CSI-sharing in the separate stream case. MMSE precoders were derived for both transmission schemes, and stream assignment strategies were proposed accordingly. Our results demonstrate that the conventional hardening bound, effective for single-antenna users, is inadequate for multi-antenna configurations; hence, we introduced a novel pilot-based downlink LMMSE channel estimation scheme along with its corresponding capacity bound. Through simulations, we demonstrated how this new approach can significantly improve the SEs and deliver performance similar to the perfect CSI case. We compared the three transmission methods and demonstrated that the same stream method outperforms both other methods. However, the separate stream transmission with CSI-sharing method performs close to the same stream transmission and it has less fronthaul requirement. Therefore, it is a feasible option. We have also showcased the impact of varying numbers of users, APs, ASD, and user and AP antennas. We have seen that even with the high number of users and user antennas, we can achieve good performance despite the increased amount of interference. Moreover, we have seen that spatial correlation has a significant impact on performance as the interference here is both among user antennas and users.
@@ A-B Derivation of the MMSE Precoder
We will derive the MMSE precoder considering a virtual uplink and use a uplink-downlink duality similar to [4, Ch. 6]. As we want to send separate streams in the downlink, we consider transmitting separate streams to the APs in the uplink. We write the received signal at AP $l$ with precoders $\boldsymbol{\Gamma}_{ji}$ as

### v2

@@ VII-B Impact of Spatial Correlation
VII-C Impact of the Number of Users
@@
VII-D Impact of the Number of Antennas per User
@@
VII-E Impact of the Coherence Block Length
In Fig. 6, we look at the impact of the coherence block length on the SE with the ZF combiner. We vary the number of user antennas as in Fig. 5 and will revisit the issue mentioned in the former subsection. We note that the number of pilots increases linearly with the number of user antennas. Having more antennas should not degrade the SE, but it does when using ZF with $\tau_{\mathrm{c}}=200$. The reason is that the SE during data transmission increases very little compared to the reduction in the pre-log factor caused by the growing pilot overhead. However, if $\tau_{\mathrm{c}}$ were higher, the extra pilot overhead would be less impactful, relatively speaking. We can see that when $\tau_{\mathrm{c}}=1000$, the SE increases with $M$ for the considered range of antennas. Hence, the benefit of having more user antennas is particularly large in low-mobility cases. From Fig. 5 and Fig. 6, we see that the MMSE combiner is a better option than the ZF combiner. We need $\tau_{\mathrm{c}}$ to be sufficiently larger if the ZF combiner is to be used.
VII-F Comparison of the Capacity Bounds
@@
Fig. 7 shows the average SE obtained with Method 1 for $M=1$ antenna per user with i.i.d. Rayleigh fading. This is the most common setup in previous works, including the textbook [demir2021foundations]. There is little difference between the curves since channel hardening occurs in this case, so one can omit downlink pilots and use the hardening bound without much performance loss.
Fig. 8 considers the same setup but with $M=2$. As discussed earlier in the paper, we cannot rely on channel hardening because only the diagonal entries of the effective channel harden, while the inter-stream interference in the off-diagonal entries does not vanish. Consequently, there is a huge gap between the perfect CSI and no CSI cases for $M>1$. However, using estimates of the downlink effective channels, as proposed in Theorem 1 leads to vastly higher performance than in the case of no CSI. The difference between the perfect CSI and the proposed method is due to two primary factors. One is the ability of the combiner filter to suppress interference and mitigate estimation error, which is inherently dependent on the effective channel estimates. Secondly, it is affected by the increased channel estimation overhead. Fig. 9 considers the case with the single AP and $N=80$ antennas, where the total number of antennas matches the cell-free massive MIMO case. We observe that the gap between the proposed method and the hardening bound remains significant in this case. Therefore, we also conclude that the gap is not due to the correlation or distribution of the APs, but rather a result of the multiple antennas on the users. These simulation results are in line with the analytical results given in Section III
In [demir2021foundations], it is briefly mentioned that multi-antenna users can be treated as multiple single-antenna users placed close to each other. Here, if we assume that each serving AP transmits a separate stream to each antenna of user $k$, our model aligns with that assumption. In Fig. 10, we compare the average SE achieved by this sub-optimal method with our proposed Method 3. The results are shown as a function of the number of antennas on each AP, and we consider both the proposed bound and the hardening bound. We observe that the hardening bound is close to the proposed bound when using the sub-optimal method. Therefore, the system can operate without downlink pilots with this transmission method but rely on channel hardening, leading to less channel estimation overhead. However, we see that the SEs achieved by this sub-optimal transmission method are significantly lower than those of Methods 1-3. This is due to the loss of beamforming gain and increased inter-stream interference. These results tell us that we should not treat a multi-antenna user as multiple single-antenna users, and we need to do pilot-based downlink channel estimation in order to have high SE with multiple antenna users. We also observe that, although the structure of the precoder is block diagonal for Method 3, the hardening bound isn’t tight. We omitted the SE results for Method 2 as these are similar to Method 3.
@@
Figure 8: Achievable per user SE as a function of the number of AP antennas with M=2M=2.444If one would double the pilot length for the perfect CSI case to model that we obtain perfect channel estimates by downlink pilots, the curve would drop a few percent.
Figure 9: Achievable per user SE as a function of the number of AP antennas with $M=2$ and a single AP.
@@
VIII Conclusions
In this work, we present a novel analysis of cell-free massive MIMO systems with multi-antenna users, where each user can receive multiple spatial streams. We analyzed the transmission of these streams from either multiple APs, which requires data sharing, or from separate APs, and compared CSI-sharing and no CSI-sharing in the separate stream case. MMSE precoders were derived for both transmission schemes, and stream assignment strategies were proposed accordingly. Our results demonstrate that the conventional hardening bound, effective for single-antenna users, is inadequate for multi-antenna configurations; hence, we introduced a novel pilot-based downlink LMMSE channel estimation scheme along with its corresponding capacity bound. Through simulations, we demonstrated how this new approach can significantly improve the SEs and deliver performance similar to the perfect CSI case. We compared the three transmission methods and demonstrated that the same stream method outperforms both other methods. However, the separate stream transmission with the CSI-sharing method performs close to the same stream transmission, and it has less fronthaul requirement. Therefore, it is a feasible option. We have also showcased the impact of varying numbers of users, APs, ASD, and user and AP antennas. We have seen that even with the high number of users and user antennas, we can achieve good performance despite the increased amount of interference. Moreover, we have seen that spatial correlation has a significant impact on performance as the interference here is both among user antennas and users.
@@ A-B Derivation of the MMSE Precoder
We will derive the MMSE precoder considering a virtual uplink and use a uplink-downlink duality similar to [demir2021foundations, Ch. 6]. As we want to send separate streams in the downlink, we consider transmitting separate streams to the APs in the uplink. We write the received signal at AP $l$ with precoders $\boldsymbol{\Gamma}_{ji}$ as
@@
[Uncaptioned image] Eren Berk Kama (Member, IEEE) received the B.S. and MSc. degrees in electrical and electronics engineering from Middle East Technical University, Ankara, Turkey, in 2019 and 2022, respectively. Since 2022, he has been a Ph.D. student at KTH Royal Institute of Technology, Stockholm, Sweden. His research interests are wireless communications and signal processing.
[Uncaptioned image] Junbeom Kim (Member, IEEE) received the B.S. and Ph.D. degrees in electronics and information engineering from Jeonbuk National University, Jeonju, Korea, in 2018 and 2023, respectively. In 2023, he was a Post-Doctoral Research Fellow at KTH Royal Institute of Technology, Stockholm, Sweden. Since 2023, he has been with Gyeongsang National University, Jinju, Korea, where he is currently an Assistant Professor with the Department of AI Information Engineering. His research interests include wireless communication technologies in PHY layer, signal processing optimization and machine learning.
[Uncaptioned image] Emil Björnson (Fellow, IEEE) received the M.S. degree in engineering mathematics from Lund University, Sweden, in 2007, and the Ph.D. degree in telecommunications from the KTH Royal Institute of Technology, Sweden, in 2011. He is now a Professor of Wireless Communication at the KTH Royal Institute of Technology, Stockholm, Sweden. He is an IEEE Fellow, Digital Futures Fellow, Wallenberg Academy Fellow, and Clarivate Highly Cited Researcher. He has a podcast and YouTube channel called Wireless Future. His research focuses on multi-antenna communications and radio resource management, using methods from communication theory, signal processing, and machine learning. He has authored four textbooks and published much simulation code. He has received the 2018 and 2022 IEEE Marconi Prize Paper Awards in Wireless Communications, the 2019 EURASIP Early Career Award, the 2019 IEEE Communications Society Fred W. Ellersick Prize, the 2019 IEEE Signal Processing Magazine Best Column Award, the 2020 Pierre-Simon Laplace Early Career Technical Achievement Award, the 2020 CTTC Early Achievement Award, the 2021 IEEE ComSoc RCC Early Achievement Award, the 2023 IEEE ComSoc Outstanding Paper Award, and the 2024 IEEE Stephen O. Rice Prize.